OpenStudy (anonymous):

what is the smallest number that is divibsile by both 45,65 plz give me explaination

OpenStudy (anonymous):

45 = 3*3*5 65 = 5*13 So 3*3*5*13 = 585

OpenStudy (anonymous):

U asked this question in another post, u should ask there not make a new post...

OpenStudy (anonymous):

45 = 3 * 3 * 5 65 = 13 * 5 Take common of both = 5 Multiply remaing factors with this ie 5 * 3* 3 * 13 = 585

OpenStudy (anonymous):

LCM (45, 65) = 585

OpenStudy (anonymous):

the factors of 45= 3*3*5 factors of 65=13*5 the factors need to be prime numbers, now to find smallest number you need to find the HCF HCF is to take the common multiple from the 2 numbers, which is 5 hence answer is 5